OpenStudy (_.mufasa._):

What is the solution of sqrt 2x+4 - sqrt x=2? A. x = 0 and x = -16 B. x = 0 and x = 16 C. x = 0 D. x = 16 and x = -16

OpenStudy (_.mufasa._):

\[\sqrt{2x+4} - \sqrt{x=2}\]

OpenStudy (anonymous):

https://mathway.com/ this website will give u the answer!! good luck!